# Table Meets LLM: Can Large Language Models Understand Structured Table Data? A Benchmark and Empirical Study

### v4

@@ Abstract.
Large language models (LLMs) are becoming attractive as few-shot reasoners to solve Natural Language (NL)-related tasks. However, there is still much to learn about how well LLMs understand structured data, such as tables. Although tables can be used as input to LLMs with serialization, there is a lack of comprehensive studies that examine whether LLMs can truly comprehend such data. In this paper, we try to understand this by designing a benchmark to evaluate the structural understanding capabilities (SUC) of LLMs. The benchmark we create includes seven tasks, each with its own unique challenges, e.g., cell lookup, row retrieval, and size detection. We perform a series of evaluations on GPT-3.5 and GPT-4. We find that performance varied depending on several input choices, including table input format, content order, role prompting, and partition marks. Drawing from the insights gained through the benchmark evaluations, we propose self-augmentation for effective structural prompting, such as critical value / range identification using internal knowledge of LLMs. When combined with carefully chosen input choices, these structural prompting methods lead to promising improvements in LLM performance on a variety of tabular tasks, e.g., TabFact($\uparrow 2.31\%$), HybridQA($\uparrow 2.13\%$), SQA($\uparrow 2.72\%$), Feverous($\uparrow 0.84\%$), and ToTTo($\uparrow 5.68\%$). We believe that our open source [^1] Please find code and data of the paper at https://github.com/microsoft/TableProvider. benchmark and proposed prompting methods can serve as a simple yet generic selection for future research.
@@ 3.2. Task Design
Table Partition
What is the first token (cell value instead of separator —) of the given table? What is the end token (cell value instead of separator —) of the given table? Answer questions one by one and use — to split the answer.
Cell Lookup
What is the position of the cell value cell_value? Use row index and column index to answer
Reverse Lookup
What is the cell value of row index, column index ? Only output the cell value without other information
Column Retrieval
What is the column name with the index column_idx of the following table? Only give the column name without any explanation
Row Retrieval
What are the cell values of the row_idx row in following table? Only list the cell values one by one using — to split the answers
Size Detection
How many rows in the table? How many columns in the table. Answer the questions one by one and use — to split the answer
Merged Cell Detection
What is the column index of the cell which span is over 1. use — to split the answer (e.g., 3 — 4), the column index starts from 0. If there’s no answer, return None
@@ 4. Structural Prompting
Formally, self-augmented prompting is a simple idea that utilizes prompting twice to leverage the capabilities of LLMs in understanding structured data, as shown in Figure 3. In the first prompt, the original task of “requesting” information is replaced with a simple demand to “Identify critical values and ranges of the last table related to the statement”. Each demand represents an important aspect of structural information. The purpose of this replacement is to unlock the reasoning abilities of LLMs for complex reasoning over structured data. The prompted text is then fed into the LLM model, which generates a subsequent sentence containing specific structural information. In the second prompt, the generated subsequent sequence is appended to the task request and fed into the LLM model to generate the final answer. Refer to Table 4 for a comparison of the experiment results using self-augmented prompting.
Table 2. Micro results of the benchmark. Change order (Xie et al., 2022) refers to put external text (like questions, statement) ahead of tables. Noted that ”GPT-4” refers to the evaluation outcomes utilizing the GPT-4 model. Given the resource-intensive nature of GPT-4 calls, we only conducting the GPT-4 inference test on a subset of 300 samples (randomly sampled) from each task set. Each column follows the roles of graded color scale, i.e., the deeper color refers to better perf.
@@
Table 3. Micro ablation results of the input designs over benchmark.
@@
GPT-4 w/ Lan. HTML 98.32% 1.65% 73.34% 29.34% 59.45% 12.12% 69.32% 5.99% 50.19% 8.19% 83.43% 16.43% 81.28% 4.61%
5. Experiments
5.1. Experiment Settings
5.2. Results
5.2.1. Benchmark Highlights
Table 4. Downstream tasks evaluation. SA. refers to Self-augmented Prompting.

### v5

@@ Abstract.
Large language models (LLMs) are becoming attractive as few-shot reasoners to solve Natural Language (NL)-related tasks. However, the understanding of their capability to process structured data like tables remains an under-explored area. While tables can be serialized as input for LLMs, there is a lack of comprehensive studies on whether LLMs genuinely comprehend this data. In this paper, we try to understand this by designing a benchmark to evaluate the structural understanding capabilities of LLMs through seven distinct tasks, e.g., cell lookup, row retrieval and size detection. Specially, we perform a series of evaluations on the recent most advanced LLM models, GPT-3.5 and GPT-4 and observe that performance varied with different input choices, including table input format, content order, role prompting, and partition marks. Drawing from the insights gained through the benchmark evaluations, we propose self-augmentation for effective structural prompting, such as critical value / range identification using internal knowledge of LLMs. When combined with carefully chosen input choices, these structural prompting methods lead to promising improvements in LLM performance on a variety of tabular tasks, e.g., TabFact($\uparrow 2.31\%$), HybridQA($\uparrow 2.13\%$), SQA($\uparrow 2.72\%$), Feverous($\uparrow 0.84\%$), and ToTTo($\uparrow 5.68\%$). We believe that our open source [^1] Please find code and data of the paper at our temporary repo at https://anonymous.4open.science/r/StructuredLLM-76F3/README.md.
Please note that we may replace the private preview with an official one in https://github.com/microsoft/TableProvider at any time.
benchmark and proposed prompting methods can serve as a simple yet generic selection for future research.
@@ 3.2. Task Design
Table Partition What is the first token (cell value instead of separator —) of the given table? What is the end token (cell value instead of separator —) of the given table? Answer questions one by one and use — to split the answer.
Cell Lookup What is the position of the cell value cell_value? Use row index and column index to answer
Reverse Lookup What is the cell value of row index, column index ? Only output the cell value without other information
Column Retrieval What is the column name with the index column_idx of the following table? Only give the column name without any explanation
Row Retrieval What are the cell values of the row_idx row in following table? Only list the cell values one by one using — to split the answers
Size Detection How many rows in the table? How many columns in the table. Answer the questions one by one and use — to split the answer
Merged Cell Detection What is the column index of the cell which span is over 1. use — to split the answer (e.g., 3 — 4), the column index starts from 0. If there’s no answer, return None
@@ 4. Structural Prompting
Table 2. Micro results of the benchmark. Change order refers to place external text such as questions or statements before tables as described in (Xie et al., 2022). ”GPT-4” refers to the task performance of GPT-4 model. Given the resource-intensive nature of GPT-4 calls, we conduct inference test on a random subset of 300 samples from each task set. Each column in the results uses a graded color scale where deeper colors signify superior performance.
@@
Table 3. Micro ablation results of the input designs over benchmark. All the $\Delta$ score denotes the performance gap compared to the first row, which includes all input design variants. Blue cells indicate an increase in performance (positive $\Delta$) while the red cells indicate an decrease in performance (negative $\Delta$). The intensity of the color corresponds to the magnitude of the change. The line marked with indicates the highest average performance (65.43%) across seven key tasks, achieved through the use of HTML markup with format explanations, role prompts, and by preserving the original sequence of inputs, i.e., placing external texts like questions and statements behind the tables. See detailed ablation results in Table 9.
@@
Self-augmented prompting is a straightforward idea that invokes a LLM twice to enhance its ability to understand structured data, as illustrated in Figure 3. This method leverages the LLM’s capacity to extract and process complex data by breaking down the task into two distinct phases. In the first stage, the model is prompted to focus on identifying crucial values and patterns in the data. Instead of directly asking the LLM to solve the primary task, it is initially tasked with extracting significant data points or determining ranges that are relevant to the downstream process. This is crucial because these extracted elements are key to building a better understanding of the context and the specific details needed for more accurate results. Once the relevant data has been identified and outlined by the LLM, the second stage involves using this refined data as part of a new prompt that now focuses on the original task. By incorporating the results from the first prompt into the second, the process effectively guides the LLM to consider these critical details while generating the final answer, enhancing both the relevance and accuracy of the response.
The merit of this approach lies in its capacity to transform a general query into a structured two-step inquiry, optimizing the model’s analytical prowess before applying it to the actual problem-solving task. This not only streamlines the reasoning process but also significantly improves the outcome by structuring the interaction with the LLM around the specific insights needed to address the query comprehensively. Refer to Table 4 for a comparison of self-augmented prompting experimental results.
Table 4. Comparison of self-augmented and 1-shot prompting across various design variants on downstream tasks. Refer to the table 6 for details on different self-augmented prompting options.
@@
Table 6. Instruction for self-augmented prompting in Section 4.
Method Instruction
Format explanation Generate short format specification and description of the table within five sentences.
Critical values and ranges identification Identify critical values and ranges of the table related within five sentences.
Structural information description Describe structural information, patterns and statistics of the table within five sentences.
5. Experiments
5.1. Experiment Settings
5.2. Results
5.2.1. Benchmark Highlights
NL+Sep vs. Markup Lan.